# Effects of particle-hole fluctuations on the superfluid transition in two-dimensional atomic Fermi gases

### v1

@@ II.2 BKT Criterion
At $T=0$ where $\mu_{\text{p}}=0$, Eqs. (1) and (2) give $\mu=E_{\text{F}}-\epsilon_{\text{B}}/2$ and $\Delta=\sqrt{2E_{\text{F}}\epsilon_{\text{B}}}$ [45]. At finite temperature, the summation in Eq. (4) analytically leads to $\frac{a_{0}}{a_{1}}-\sqrt{\frac{a_{0}}{a_{1}}(\frac{a_{0}}{a_{1}}-4\mu_{\text{p}})}=2T\ln(1-e^{-\mathcal{D}_{\text{B}}})$, which reduces to $\mu_{\text{p}}=T\ln(1-e^{-\mathcal{D}_{\text{B}}})$ in the BEC regime. Here $n_{\text{B}}=a_{0}\Delta^{2}$ represents the pair density, and $\mathcal{D}_{\text{B}}$ is the bosonic phase space density. Thus $\mu_{\text{p}}$ is primarily determined by $\mathcal{D}_{\text{B}}$ at given temperature, indicating that the BKT transition may occur when $\mathcal{D}_{\text{B}}$ is large enough so that $\mu_{\text{p}}$ becomes sufficiently close to zero [46, 17, 24, 18]. Moreover, in the BEC regime, a large $\mathcal{D}_{\text{B}}$ provides a sharp peak distribution of the bosonic number density at zero momentum via $b(-\mu_{\text{p}})=e^{\mathcal{D}_{\text{B}}}-1$, signaling a quasi-condensation [24]. Alternatively, when the phase space density becomes large enough, the wave functions of neighboring pairs start to overlap with each other, and hence help to establish quasi-long-range phase coherence.
When approached from the high-temperature region [47, 4], the bosonic BKT transition occurs when $\mathcal{D}_{\text{B}}(T)$ reaches a critical value $\mathcal{D}_{\text{B}}(T_{\text{BKT}})$ [46]. Estimates of $\mathcal{D}_{\text{B}}(T_{\text{BKT}})$ for fermionic superfluids are provided in Refs. [16, 17], with values ranging from approximately 4.9 to 6.45. In comparison, the analogous systems in atomic Bose gases typically hover around 8 [6]. The lowest value, $\mathcal{D}_{\text{B}}(T_{\text{BKT}})=4.9$, which is the closest to the factor of 4 in the usual BKT relation, was reported to offer the best fit for the experimental results on Fermi gases [17]. Thus we choose $\mathcal{D}_{\text{B}}(T_{\text{BKT}})=4.9$, and the BKT transition temperature $T_{\text{BKT}}$ is determined by
$\frac{n_{\text{B}}}{M_{\text{B}}}=\frac{4.9}{2\pi}T_{\text{BKT}}\,.$ (5)
@@ II.3 Contributions of the Particle-Hole Channel
Following Ref. [32], we introduce the contribution of the particle-hole channel, by renormalizing the pairing strength in $t(Q)$, which leads to a new full $T$-matrix $t_{\text{2}}(Q)$ that includes both particle-particle and particle-hole contributions. The expression for $t_{\text{2}}(Q)$ is given by
@@
which self-consistently includes the self-energy feedback, with $K,K^{\prime}$ being the external fermion momentum. Here the particle-hole channel $T$-matrix $t^{-1}_{\text{ph}}(Q^{\prime})=U^{-1}+\chi_{\text{ph}}(Q^{\prime})$ describes the particle-hole scattering, and the particle-hole susceptibility $\chi_{\text{ph}}(Q^{\prime})=\sum_{K}G(K)G_{0}(K-Q^{\prime})$, where the particle-hole momentum $Q^{\prime}=K+K^{\prime}-Q$. Moreover, assuming that the fermions near the Fermi surface dominate the particle-hole channel contributions, we replace the particle-hole susceptibility with an average $\langle{\chi_{\text{ph}}\rangle}$ on or near the Fermi surface, where the frequency part of the particle-hole susceptibility is set to 0 with $\mathrm{i}\Omega^{\prime}_{n}=0$ [48], leading to a zero imaginary part of $\chi_{\text{ph}}(Q^{\prime})$ and therefore a purely real $\langle\chi_{\text{ph}}\rangle$ [32].
We have proposed two methods for averaging $\chi_{\text{ph}}(Q^{\prime})$, referred to as level 1 and level 2, respectively. The level 1 average involves an on-shell and elastic scattering on the Fermi surface, with $|\mathbf{k}|=|\mathbf{k}^{\prime}|=k_{\mu}=\sqrt{2m\max(\mu,0)}$, where the momentum part of $\chi_{\text{ph}}(Q^{\prime})$ is determined by $|{\mathbf{q}^{\prime}}|=\left|\mathbf{k}+\mathbf{k}^{\prime}\right|=k_{\mu}\sqrt{2(1+\cos\theta)}$. Here $\chi_{\text{ph}}(Q^{\prime})$ is averaged over scattering angles $\theta$ (between $\mathbf{k}$ and $\mathbf{k}^{\prime}$). This averaging process, focusing solely on the Fermi surface, is commonly used in the literature on the studies of induced interactions. In contrast, the level 2 average considers that the states within the energy range $\xi_{\mathbf{k}}\in[-\min(\Delta,\mu),\Delta]$ of a typical s-wave superconductor are most significantly affected by pairing (for $\mu>0$). Hence, for the level 2 average, while keeping the on-shell and elastic scattering, the average is performed over a range of $|\mathbf{k}|$ such that the quasi-particle energy $E_{\mathbf{k}}\in\left[\min(E_{\mathbf{k}}),\min(\sqrt{E^{2}_{\mathbf{k}}+\Delta^{2}})\right]$, where $\min(E_{\mathbf{k}})=\Delta$ if $\mu>0$, or $\min(E_{\mathbf{k}})=\sqrt{\mu^{2}+\Delta^{2}}$ if $\mu<0$.
@@
Throughout the BCS-BEC crossover, $\langle\chi_{\text{ph}}\rangle$ remains negative as a function of the coupling strength. From Eq. (7), the particle-hole channel constitutes a renormalization of the pairing interaction, with a net effect given by replacing $1/U$ with $1/U_{\text{eff}}\equiv 1/U+\langle\chi_{\text{ph}}\rangle$. Diagrammatically, this amounts to replacing the bare interaction $U$ with the full particle-hole $T$-matrix $t_{\text{ph}}$ in the particle-particle scattering diagrams [32]. Given the negative sign of $\langle\chi_{\text{ph}}\rangle$, one can see immediately that $|U_{\text{eff}}|<|U|$. Thus $U_{\text{eff}}$ represents a weaker, screened pairing interaction.
@@ III.1 Behaviors of the particle-hole susceptibility
First, we present in Fig. 1 the (level 1) angular average of the particle-hole susceptibility at zero frequency as a function of momentum $k$, under the on-shell condition $|\mathbf{k}|=|\mathbf{k}^{\prime}|=k$. Here we focus on the unitary case at $T=T_{\text{BKT}}$ (black solid line) and zero-temperature (red dashed line), and $\langle\chi_{\text{ph}}(0,|\mathbf{k}+\mathbf{k}^{\prime}|)\rangle$ is calculated using the corresponding solution of $(\Delta,\mu,\mu_{\text{p}})$ at $T_{\text{BKT}}/T_{\text{F}}=0.079$ and $\ln(k_{\text{F}}a_{\text{2D}})=0$, solved in the absence of particle-hole fluctuations. The slight difference between these two curves reveals a weak temperature dependence. Importantly, both curves show a strong momentum dependency, with the amplitude decreasing monotonically as the momentum rises. Note that the momentum dependencies in 2D appear distinct from those in 3D [32], where $\langle\chi_{\text{ph}}(0,|\mathbf{k}+\mathbf{k}^{\prime}|)\rangle$ exhibits a nonmonotonic momentum dependence at low $T$ for the unitary case.
@@
Shown in Fig. 2 is $\langle\chi_{\text{ph}}\rangle$ at $T=T_{\text{BKT}}$ as a function of $\ln(k_{\text{F}}a_{\text{2D}})$ throughout the BCS-BEC crossover, averaged at both level 1 (red dashed) and level 2 (black solid curve), along with its BCS limit $1/4\pi\approx 0.0796$, given by $\langle\chi_{\text{ph}}\rangle=-{m}/{2\pi}$. The value of $\langle\chi_{\text{ph}}\rangle/2m$ decreases monotonically with increasing interaction strength and exhibits an exponential behavior in the deep BEC regime for $\ln(k_{\text{F}}a_{\text{2D}})<-2$. As the gap diminishes, the average at both levels converges in the BCS limit. However, in the crossover regime, the two levels differ significantly, with a smaller absolute value for the level 2 average. This can be understood from Fig. 1; compared with level 1 averaging, level 2 averaging involves a range of $k$’s so that the larger $k$ contributions dominates due to its larger phase space area, leading to a smaller absolute value of the average. In other words, level 1 averaging on the Fermi surface only leads to a significant over-estimate of the particle-hole contributions in the unitary regime.
The asymptotic behavior of $\langle\chi_{\text{ph}}\rangle$ can be readily solved analytically in both the BCS and the BEC limits. In the weak coupling limit, where $\ln(k_{\text{F}}a_{\text{2D}})\rightarrow\infty$, $\Delta\rightarrow 0$, and $T\leq T_{\text{BKT}}\rightarrow 0$, the two levels of average converge to $\chi_{\text{ph}}(Q^{\prime})\approx\sum_{K}G_{0}(K)G_{0}(K-Q^{\prime})$. Under the on-shell condition $\xi_{\mathbf{k}}=\xi_{\mathbf{k}-\mathbf{q}^{\prime}}$, the integrand becomes the derivative of the Fermi function and one readily obtains
@@
We emphasize that this result is independent of the density, due to the constant density of states in 2D. This is to be contrasted with its 3D counterpart, which is proportional to the on-shell momentum $k$. In the strong coupling limit, where $\ln(k_{\text{F}}a_{\text{2D}})\rightarrow$-$\infty$, we have $|\mu|\gg\Delta\gg\epsilon_{\text{F}}$, which indicates that $E_{\mathbf{k}}\approx\xi_{\mathbf{k}}\approx|\mu|$. Thus particle-hole fluctuations are exponentially suppressed, so that $\langle\chi_{\text{ph}}\rangle$ approaches zero.
III.2 Effect of particle-hole channel on the BKT transition
@@
In this section, we explore the impact of the particle-hole channel on the pairing phenomena and the BKT transition of a 2D Fermi gas. We begin by examining the behavior of the pairing gap $\Delta$ at zero temperature with and without the particle-hole channel contributions. Plotted in Fig. 3(a) is $\Delta$ as a function of interaction. For comparison, we plot the results both without (black solid line) and with the particle-hole channel effect, with the particle-hole susceptibility $\langle{\chi_{\text{ph}}\rangle}$ averaged at level 1 (red dashed curve) and level 2 (blue dot-dashed curve), respectively. The corresponding $\langle{\chi_{\text{ph}}\rangle}$ is shown in Fig. 3(b). For all cases, $\Delta$ decreases monotonically from BEC to BCS, as expected. A substantial reduction of $\Delta$ by the particle-hole fluctuations occurs in the crossover and BCS regimes. In the BEC regime, $\langle{\chi_{\text{ph}}\rangle}$ approaches zero, rendering the particle-hole channel effect negligible. Consistent with Fig. 3(b), level 2 averaging showed a weaker particle-hole effect, and thus a smaller reduction of $\Delta(T=0)$. Note that in Fig. 3(b), there exists a kink around $\mu=0$ in $\langle{\chi_{\text{ph}}\rangle}$ for both levels of averaging, mainly because the Fermi surface shrinks to zero abruptly with a finite slope as a function of increasing pairing strength. The Fermi function in the integrand of $\langle{\chi_{\text{ph}}\rangle}$ becomes a step function at zero $T$, resulting in a delta function in the derivative of the integrand of $\langle{\chi_{\text{ph}}\rangle}$ and hence a discontinuity when crossing $\mu=0$ (See Appendix A for details). This slope discontinuity becomes more prominent in the level 2 average, since the range for $k<k_{\mu}$ in the average shrinks to zero abruptly when $\mu=0$  [^1] It should be noted that $\mu=0$ occurs at $\ln(k_{F}a_{\text{2D}})\approx-0.25$ and -0.175 for level 1 and 2 averaging, respectively, as they are calculated with the self-consistent solutions of $T_{\text{BKT}}$..
Figure 4: Effect of the particle-hole channel contributions on (a) $T_{\text{BKT}}$, (b) $\Delta$, (c) $\mu$, (d) $\mu_{\text{p}}$, (e) $n_{\text{B}}$ and (f) $M_{\text{B}}$ versus $\ln(k_{\text{F}}a_{\text{2D}})$, with the average $\langle\chi_{\text{ph}}\rangle/2m$ calculated at level 1 (red dashed) and level 2 (blue dot-dashed curves), respectively. They should be compared with the results without the particle-hole effect (black solid curves).
For the effect of the particle-hole channel on the behavior of the BKT transition temperature $T_{\text{BKT}}$, we first analytically estimate the ratio between the two BKT transition temperatures in the BCS limit with and without the particle-hole channel at the same coupling strength, where the latter is denoted as $T_{\text{BKT}}^{\text{BCS}}$. Here $\Delta\rightarrow 0$, so that the following summation can be performed analytically,
@@
where $\gamma\approx 0.5772157$ is Euler’s constant. (A detailed derivation can be found in Appendix B.) Substituting the above relation for the corresponding term in Eqs. (1) and (7), we obtain
@@
Next, we present in Fig. 4 the effect of the particle-hole channel on the evolution of (a) $T_{\text{BKT}}$, along with (b) $\Delta$, (c) $\mu$, (d) $\mu_{\text{p}}$, (e) $n_{\text{B}}$, and (f) $M_{\text{B}}$ at $T_{\text{BKT}}$ as a function of $\ln(k_{\text{F}}a_{\text{2D}})$. The results without particle-hole fluctuations are presented as the black solid curves. Starting from the weak coupling BCS limit (black curve), as the interaction strength increases, $T_{\text{BKT}}$ increases, and then reaches a maximum in the intermediate regime, where $\mu_{\text{p}}$ reaches a minimum simultaneously. As the interaction strength increases further past the maximum, $T_{\text{BKT}}$ decreases and reaches a minimum near unitarity $\ln(k_{\text{F}}a_{\text{2D}})=0$, where $\mu=0$. Beyond this point, the system enters the BEC regime, where all fermions pair up with $2n_{\text{B}}/n\approx 1$. The behavior of $T_{\text{BKT}}$ is then influenced by the shrinking pair size, as indicated by a decrease in the pair mass $M_{\text{B}}$ toward $2m$, reaching a BEC asymptote $T_{\text{BKT}}/T_{\text{F}}\approx 0.109$. The gap $\Delta$ consistently increases with pairing strength. In comparison, the particle-hole channel contributions cause a non-uniform shift of all curves toward the BEC regime on the right. This shift is the largest in the BCS limit, and vanishes in the BEC regime, as indicated in Fig. 2. It exhibits a strong dependence on $\ln(k_{\text{F}}a_{\text{2D}})$ in the crossover regime. Now the maximum of $T_{\text{BKT}}$ occurs closer to unitarity, and the locate for $\mu=0$ is shifted into the BEC regime. All $T_{\text{BKT}}$ curves with and without the particle-hole effect converge to the same BEC asymptote. From Fig. 4(c), we have a tiny $|\mu_{\text{p}}|\leq 1.2\times 10^{-3}E_{\text{F}}$ throughout the BCS-BEC crossover for all cases. This ensures that $t(Q)$ remains highly peaked at $Q=0$ and thus validates our pseudogap approximation for the self-energy, $\Sigma(K)\approx-\Delta^{2}G_{0}(-K)$.
@@
Figure 5: Effect of the particle-hole channel contributions on behaviors of $\Delta$ (top row) and $\mu_{\text{p}}$ (bottom row), as a function of $T/T_{\text{BKT}}$, with $\ln(k_{\text{F}}a_{\text{2D}})=-1,0,2$ from left to right for the BCS, unitary, and BEC regimes, respectively. The black solid curve represents calculations without the particle-hole channel, while the red dashed and green dot-dashed curves include the particle-hole channel effect, using $\langle\chi_{\text{ph}}\rangle/2m$ under level 1 and level 2 averaging, respectively.
@@
Now we investigate in Fig. 5 the evolution of $\Delta$ (top row) and $\mu_{\text{p}}$ (bottom row) as a function of reduced temperature $T/T_{\text{BKT}}$ in the BCS, unitary, and BEC regimes from left to right without (black) and with (red and green) the particle-hole channel effect. The gap $\Delta$ remains nearly constant except in the BCS case (f), where a significant decrease can be discerned near $T_{\text{BKT}}$. The value of $\Delta$ with the particle-hole channel effect is reduced from the counterpart without the particle-hole channel effect by a factor of $1/e$ in the BCS limit. This reduction factor becomes smaller in the unitary and BEC regimes, consistent with the zero $T$ and $T_{\text{BKT}}$ gap behaviors as shown in Fig. 3(a) and Fig. 4(b). At the same time, for all cases, $\mu_{\text{p}}$ decreases continuously to zero as $T$ decreases, consistent with $\mu_{\text{p}}=0$ at $T=0$ for a true long-range-order ground state. Note that below $0.6T_{\text{BKT}}$, $\mu_{\text{p}}$ comes essentially zero.

### v2

@@ II.2 BKT Criterion
The Nelson-Kosterlitz condition (NKC) provides a criterion for determining $T_{\text{BKT}}$. However, there are issues with this condition. First of all, due to the Galilean translational invariance, one has the zero $T$ superfluid density $n_{s}(T=0)=n$, independent of the pairing strength. Furthermore, from the fermionic perspective, the fermion mass $m$ is a constant. Hence it is hard to imagine that $n_{s}(T_{\text{BKT}})/m$ would be able to properly capture the interaction effect. While it is often assumed $M_{\text{B}}=2m$ in a NKC-based calculation, however, the correct $M_{B}$ should necessarily be renormalized by the interactions. Therefore, it is reasonable to suspect that NKC is not quantitatively reliable for a fermionic system.
Here we choose the criterion of critical phase space density ${\cal D}_{B}$ to determine $T_{\text{BKT}}$, as established by Prokof’ev et al via quantum Monte Carlo (QMC) simulations of a Bose gas [41]. When approached from high temperature [30, 22], the bosonic BKT transition occurs when $\mathcal{D}_{\text{B}}(T)$ reaches a critical value $\mathcal{D}_{\text{B}}(T_{\text{BKT}})$ such that the boson wavefunctions start to overlap with each other [42]. Hence we have
$\frac{n_{\text{B}}}{M_{\text{B}}}=\frac{{\cal D}_{\text{B}}}{2\pi}T_{\text{BKT}}\,$, (5)
where $n_{\text{B}}$ and $M_{\text{B}}$ are determined as described above. In the deep BEC regime, we have $n_{\text{B}}=n/2$ and $M_{\text{B}}=2m$, and hence the asymptotic behavior $T_{\text{BKT}}=T_{\text{F}}/2{\cal D}_{\text{B}}$. This would reduce to the NKC when ${\cal D}_{\text{B}}=4$.
[…]
$\tilde{g}=\frac{5.03085}{1.23629-\ln(k_{\text{F}}a_{\text{2D}})}.$ (6)
See App. B for details. The corresponding ${\cal D}_{\text{B}}$ is given by the blue curve in Fig. 7. The fit also reveals that there exists a quantitative difference between $a_{\text{2D}}$ and either the potential radius or the pair size.
[…]
At $T=0$ where $\mu_{\text{p}}=0$, Eqs. (1) and (2) give $\mu=E_{\text{F}}-\epsilon_{\text{B}}/2$ and $\Delta=\sqrt{2E_{\text{F}}\epsilon_{\text{B}}}$ [43]. At finite temperature, the summation in Eq. (4) analytically leads to $\frac{a_{0}}{a_{1}}-\sqrt{\frac{a_{0}}{a_{1}}(\frac{a_{0}}{a_{1}}-4\mu_{\text{p}})}=2T\ln(1-e^{-\mathcal{D}_{\text{B}}})$, which reduces to $\mu_{\text{p}}=T\ln(1-e^{-\mathcal{D}_{\text{B}}})$ in the BEC regime. Here use has been made of the relation $\mathcal{D}_{\text{B}}=2\pi n_{\text{B}}/M_{\text{B}}T$. Thus $\mu_{\text{p}}$ is primarily determined by $\mathcal{D}_{\text{B}}$ at given temperature, indicating that when the BKT transition occurs $\mathcal{D}_{\text{B}}$ is large enough so that $\mu_{\text{p}}$ becomes sufficiently close to zero [42, 38, 50, 49]. Moreover, in the BEC regime, a large $\mathcal{D}_{\text{B}}$ provides a sharp peak distribution of the bosonic number density at zero momentum via $b(-\mu_{\text{p}})=e^{\mathcal{D}_{\text{B}}}-1$, signaling a quasi-condensation [50]. Alternatively, when ${\cal D}_{\text{B}}(T)$ becomes large enough, the wave functions of neighboring pairs start to overlap with each other, and hence help to establish quasi-long-range phase coherence.
@@ II.3 Contributions of the Particle-Hole Channel
Following Ref. [17], we introduce the contribution of the particle-hole channel, by renormalizing the pairing strength in $t(Q)$, which leads to a new full $T$-matrix $t_{\text{2}}(Q)$ that includes both particle-particle and particle-hole contributions. The expression for $t_{\text{2}}(Q)$ is formally given by
@@
which self-consistently includes the self-energy feedback, with $K,K^{\prime}$ being the external fermion momentum. Here the particle-hole channel $T$ matrix $t^{-1}_{\text{ph}}(Q^{\prime})=U^{-1}+\chi_{\text{ph}}(Q^{\prime})$ describes the particle-hole scattering, and the particle-hole susceptibility $\chi_{\text{ph}}(Q^{\prime})=\sum_{K}G(K)G_{0}(K-Q^{\prime})$, where the particle-hole momentum $Q^{\prime}=K+K^{\prime}-Q$. Moreover, assuming that the fermions near the Fermi surface dominate the particle-hole channel contributions, we replace the particle-hole susceptibility with an average $\langle{\chi_{\text{ph}}\rangle}$ on or near the Fermi surface, where the frequency part of the particle-hole susceptibility is set to 0 with $\mathrm{i}\Omega^{\prime}_{n}=0$ [21], leading to a zero imaginary part of $\chi_{\text{ph}}(Q^{\prime})$ and therefore a purely real $\langle\chi_{\text{ph}}\rangle$ [17].
We have proposed two methods for averaging $\chi_{\text{ph}}(Q^{\prime})$, referred to as level 1 and level 2, respectively. The level 1 average involves an on-shell and elastic scattering on the Fermi surface, with $|\mathbf{k}|=|\mathbf{k}^{\prime}|=k_{\mu}=\sqrt{2m\max(\mu,0)}$, where the momentum part of $\chi_{\text{ph}}(Q^{\prime})$ is determined by $|{\mathbf{q}^{\prime}}|=\left|\mathbf{k}+\mathbf{k}^{\prime}\right|=k_{\mu}\sqrt{2(1+\cos\theta)}$. Here $\chi_{\text{ph}}(Q^{\prime})$ is averaged over scattering angles $\theta$ (between $\mathbf{k}$ and $\mathbf{k}^{\prime}$). This averaging process, focusing solely on the Fermi surface, is commonly used in the literature on the studies of induced interactions [51]. In contrast, the level 2 average considers that the states within the energy range $\xi_{\mathbf{k}}\in[-\min(\Delta,\mu),\Delta]$ of a typical s-wave superconductor are most significantly affected by pairing (for $\mu>0$). Hence, for the level 2 average, while keeping the on-shell and elastic scattering assumption, the average is performed over a range of $|\mathbf{k}|$ such that the quasi-particle energy $E_{\mathbf{k}}\in\left[\min(E_{\mathbf{k}}),\min(\sqrt{E^{2}_{\mathbf{k}}+\Delta^{2}})\right]$, where $\min(E_{\mathbf{k}})=\Delta$ if $\mu>0$, or $\min(E_{\mathbf{k}})=\sqrt{\mu^{2}+\Delta^{2}}$ if $\mu<0$.
@@
Throughout the BCS-BEC crossover, $\langle\chi_{\text{ph}}\rangle$ remains negative as a function of the coupling strength. From Eq. (8), the particle-hole channel constitutes a renormalization of the pairing interaction, with a net effect given by replacing $1/U$ with $1/U_{\text{eff}}\equiv 1/U+\langle\chi_{\text{ph}}\rangle$. Diagrammatically, this amounts to replacing the bare interaction $U$ with the full particle-hole $T$-matrix $t_{\text{ph}}$ in the particle-particle scattering diagrams [17]. Since $\langle\chi_{\text{ph}}\rangle<0$, one has $|U_{\text{eff}}|<|U|$. Thus $U_{\text{eff}}$ represents a weaker, screened pairing interaction.
@@ III.1 Behaviors of the particle-hole susceptibility
First, we present in Fig. 1 the (level 1) angular average of the particle-hole susceptibility at zero frequency as a function of momentum $k$, under the on-shell condition $|\mathbf{k}|=|\mathbf{k}^{\prime}|=k$. Here we focus on the unitary case at $T=T_{\text{BKT}}$ (black solid line) and zero-temperature (red dashed line), and $\langle\chi_{\text{ph}}(0,|\mathbf{k}+\mathbf{k}^{\prime}|)\rangle$ is calculated using the corresponding solution of $(\Delta,\mu,\mu_{\text{p}})$ at $T_{\text{BKT}}/T_{\text{F}}=0.079$ and $\ln(k_{\text{F}}a_{\text{2D}})=0$, solved in the absence of particle-hole fluctuations. The slight difference between these two curves reveals a weak temperature dependence. Importantly, both curves show a significant momentum dependence, with the amplitude decreasing monotonically as the momentum rises. Note that the momentum dependencies in 2D appear distinct from those in 3D [17], where $\langle\chi_{\text{ph}}(0,|\mathbf{k}+\mathbf{k}^{\prime}|)\rangle$ exhibits a nonmonotonic momentum dependence at low $T$ for the unitary case.
@@
Shown in Fig. 2 is $\langle\chi_{\text{ph}}\rangle$ at $T=T_{\text{BKT}}$ as a function of $\ln(k_{\text{F}}a_{\text{2D}})$ throughout the BCS-BEC crossover, averaged at both level 1 (red dashed) and level 2 (black solid curve), along with its BCS limit $1/4\pi\approx 0.0796$, given by $\langle\chi_{\text{ph}}\rangle=-{m}/{2\pi}$. The value of $\langle\chi_{\text{ph}}\rangle/2m$ decreases monotonically with increasing interaction strength and exhibits an exponential behavior in the deep BEC regime for $\ln(k_{\text{F}}a_{\text{2D}})<-2$. As the gap diminishes, the average at both levels converges in the BCS limit. However, in the crossover regime, the two levels differ significantly, with a smaller absolute value for the level 2 average. This can be understood from Fig. 1; compared with level 1 averaging, level 2 averaging involves a range of $k$’s so that the larger $k$ contributions dominate due to its larger phase space area, leading to a smaller absolute value of the average. In other words, level 1 averaging on the Fermi surface only leads to a significant over-estimate of the particle-hole contributions in the unitary regime.
The asymptotic behavior of $\langle\chi_{\text{ph}}\rangle$ can be readily solved analytically in both the BCS and the BEC limits. In the weak coupling limit, where $\ln(k_{\text{F}}a_{\text{2D}})\rightarrow\infty$, $\Delta\rightarrow 0$, and $T_{\text{BKT}}\rightarrow 0$, the two levels of average converge to $\chi_{\text{ph}}(Q^{\prime})\approx\sum_{K}G_{0}(K)G_{0}(K-Q^{\prime})$. Under the on-shell condition $\xi_{\mathbf{k}}=\xi_{\mathbf{k}-\mathbf{q}^{\prime}}$, the integrand becomes the derivative of the Fermi function and one readily obtains
@@
We emphasize that this result is independent of the density, due to the constant density of states in 2D. This is to be contrasted with its 3D counterpart, which is proportional to the Fermi momentum $k_{\text{F}}$. In the strong coupling limit, where $\ln(k_{\text{F}}a_{\text{2D}})\rightarrow$-$\infty$, we have $|\mu|\gg\Delta\gg E_{\text{F}}$, so that $E_{\mathbf{k}}\approx\xi_{\mathbf{k}}\approx|\mu|$. Thus particle-hole fluctuations are exponentially suppressed, so that $\langle\chi_{\text{ph}}\rangle$ approaches zero.
III.2 Effect of the particle-hole channel on BKT transition
@@
Figure 4: Effect of the particle-hole channel on (a) $T_{\text{BKT}}$, (b) $\Delta$, (c) $\mu$, (d) $\mu_{\text{p}}$, (e) $n_{\text{B}}$ and (f) $M_{\text{B}}$ as a function of $\ln(k_{\text{F}}a_{\text{2D}})$, with the average $\langle\chi_{\text{ph}}\rangle/2m$ calculated at level 1 (red dashed) and level 2 (blue dot-dashed curves), respectively. They should be compared with the results without the particle-hole contributions (black solid curves). Shown in the inset of (b) is $T_{\text{BKT}}$ over a broader range. The difference between levels 1 and 2 is not dramatic.
[…]
For the effect of the particle-hole channel on the behavior of the BKT transition temperature $T_{\text{BKT}}$, we first analytically estimate the ratio between the two BKT transition temperatures in the BCS limit with and without the particle-hole channel at the same coupling strength, where the latter is denoted as $T_{\text{BKT}}^{\text{BCS}}$. Now that $\Delta\rightarrow 0$, we obtain
@@
where $\gamma\approx 0.5772157$ is Euler’s constant. Substituting the above relation for the corresponding term in Eqs. (1) and (8), we obtain
@@
Next, we present in Fig. 4 the effect of the particle-hole channel on the evolution of (a) $T_{\text{BKT}}$, along with (b) $\Delta$, (c) $\mu$, (d) $\mu_{\text{p}}$, (e) $n_{\text{B}}$, and (f) $M_{\text{B}}$ at $T_{\text{BKT}}$ as a function of $\ln(k_{\text{F}}a_{\text{2D}})$. Shown in the inset is $T_{\text{BKT}}$ over a broad range of $\ln(k_{\text{F}}a_{\text{2D}})$. For comparison, the results without particle-hole fluctuations are presented as the black solid curves. Consider the case without particle-hole effects first. Starting from the weak coupling BCS limit (black curve), as the interaction strength increases, $T_{\text{BKT}}$ increases, and then reaches a maximum around $\ln(k_{\text{F}}a_{\text{2D}})=0.7$ in the intermediate regime, where $\mu_{\text{p}}$ reaches a minimum simultaneously. The fact there appears a maximum reflects that $n_{\text{B}}$ rises faster than $M_{\text{B}}$ in this regime (see (e,f)). As the interaction strength increases further past the maximum, $T_{\text{BKT}}$ decreases and reaches a minimum near unitarity $\ln(k_{\text{F}}a_{\text{2D}})=0$, where $\mu=0$. Beyond this point, the system enters the BEC regime, where all fermions pair up with $2n_{\text{B}}/n\approx 1$, as shown in Fig. 4(e). The behavior of $T_{\text{BKT}}$ is then dominated by the shrinking pair size, as indicated by a decrease in the pair mass $M_{\text{B}}$ toward $2m$ (Fig. 4(f)), reaching the BEC asymptotic behavior $T_{\text{BKT}}/T_{\text{F}}=1/2{\cal D}_{\text{B}}$ governed by Eq. (5). This decreasing $M_{\text{B}}$ leads to an increase and hence a minimum in $T_{\text{BKT}}$. The gap $\Delta$ in Fig. 4(b) increases consistently with pairing strength. In comparison, the particle-hole contributions cause a non-uniform shift of all curves toward the BEC regime on the right. This shift is the largest in the BCS limit, and vanishes in the BEC regime, as indicated in Fig. 2. It exhibits a strong dependence on $\ln(k_{\text{F}}a_{\text{2D}})$ in the crossover regime. Now the maximum of $T_{\text{BKT}}$ occurs closer to unitarity, and the location for $\mu=0$ is shifted into the previous BEC regime. All $T_{\text{BKT}}$ curves with and without the particle-hole effect converge to the same BEC asymptotic curve, since the particle-hole contributions vanish gradually in the BEC regime. From Fig. 4(d), we see a tiny $|\mu_{\text{p}}|\leq 1.2\times 10^{-3}E_{\text{F}}$ throughout the BCS-BEC crossover for all cases. Especially, in the BEC regime, it is uniquely determined by ${\cal D}_{\text{B}}$. This ensures that $t(Q)$ remains highly peaked at $Q=0$ and thus validates our pseudogap approximation for the self-energy.
@@
Figure 5: Effect of the particle-hole channel contributions on behaviors of $\Delta$ (top row) and $\mu_{\text{p}}$ (bottom row), as a function of $T/T_{\text{BKT}}$, with $\ln(k_{\text{F}}a_{\text{2D}})=2,0,-1$ from left to right for the BCS, unitary, and BEC regimes, respectively. The black solid curves represent calculations without the particle-hole channel, while the red dashed and green dot-dashed curves include the particle-hole channel effect, using $\langle\chi_{\text{ph}}\rangle/2m$ under level 1 and level 2 averaging, respectively.
@@
Now we investigate in Fig. 5 the evolution of $\Delta$ (top row) and $\mu_{\text{p}}$ (bottom row) as a function of reduced temperature $T/T_{\text{BKT}}$ in the BCS, unitary, and BEC regimes from left to right without (black) and with (red and green) the particle-hole channel effect. The gap $\Delta$ remains nearly constant except in the BCS case (a), where a significant decrease can be discerned near $T_{\text{BKT}}$. The pseudogap $\Delta$ with the particle-hole channel effect is reduced from the counterpart without the particle-hole effect by a factor of $1/e$ in the BCS limit. This reduction factor becomes smaller in the unitary and BEC regimes, consistent with the zero $T$ and $T_{\text{BKT}}$ gap behaviors shown in Fig. 3(a) and Fig. 4(b). At the same time, for all cases, $\mu_{\text{p}}$ increases continuously to zero as $T$ decreases, consistent with $\mu_{\text{p}}=0$ at $T=0$ for a true long-range-ordered ground state. Note that below $0.6T_{\text{BKT}}$, $\mu_{\text{p}}$ comes essentially zero. Interestingly, $|\mu_{\text{p}}|$ decreases much faster with decreasing $T$ in the BCS regime than in the unitary and BEC regimes. Conversely, should the calculation be extended to above $T_{\text{BKT}}$, $|\mu_{\text{p}}|$ would rise rapidly with $T$ so that pairs would quickly become irrelevant in this regime.